# CycleChemist: A Dual-Pronged Machine Learning Framework for Organic Photovoltaic Discovery

### v1

@@ Introduction
Current machine learning models used for predicting the power conversion efficiency (PCE) of organic photovoltaic (OPV) materials remain relatively simplistic. Limitations in both algorithm design and dataset sizes have resulted in insufficient predictive accuracy and generalizability. For example, a comprehensive survey of various embedding methods and machine learning models was conducted using 558 data points, with tree-based algorithms achieving the highest predictive correlation (approximately 0.61) (Beyond2024review). However, such a small dataset may not adequately represent mainstream OPV material systems, thereby significantly affecting generalizability. In another study, a larger dataset was compiled and a random forest model was trained, reporting an $R^{2}$ value of 0.71 (Saeki2021). Nonetheless, the reliance on experimental data as model input renders this approach impractical for large-scale screening studies. In (1060), training was performed on 1,060 donor–acceptor pairs using various graph neural network algorithms, with the Graph Attention Network (GAT) achieving the best result (reported $r$ value of 0.74). Additionally, the model presented in (300) was trained on 300 donor–acceptor pairs and involved detailed quantum mechanical calculations for each molecule during dataset preparation, introducing new quantum mechanical descriptors. On this small dataset, the gradient boosting (GB) model achieved the best performance with a reported $r$ value of 0.92. However, the complexity of the quantum mechanical (QM) calculations limits the model’s potential for large-scale screening. Similarly, another study utilizing only 100 data points incorporated numerous QM descriptors and reported an $r$ value of 0.918 (100). In summary, these methods are either trained on relatively small datasets or require complex QM calculations or additional experimental parameters, all of which limit their applicability for large-scale screening and may lead to performance that is somewhat contingent on chance.
@@
Previous research has predominantly focused on predicting the properties of photovoltaic materials, and in recent years, some scholars have begun to explore donor–acceptor design based on machine learning. According to the literature, methods for exploring high-efficiency OPV materials mainly include strategies such as disassembling and recombining existing molecular structures, as well as employing variational autoencoders (VAE) and genetic algorithms (GA) (Hutchison2023; Sun2024; ZhangSA2025). These approaches largely rely on optimizing known molecules, thereby limiting their capacity to generate and explore entirely novel molecules. Some studies do not involve molecular design at all but instead exhaustively enumerate possible donor–acceptor combinations, exploring potential high-efficiency pairs within a predefined design space (100). Moreover, most related studies employ datasets comprising only 100 to 1,500 samples (Saeki2021; Min2020; Hutchison2023), covering relatively narrow material systems, which consequently restricts the diversity of molecular fragments that the models can explore.
@@
Based on previous works (Saeki2021; Min2020), we integrated and curated multiple organic photovoltaic datasets to construct the Organic Photovoltaic Donor-Acceptor Dataset (OPV2D), comprising approximately 2,000 donor–acceptor pairs—one of the largest OPV datasets to date. Molecules are represented using SMILES (Simplified Molecular Input Line Entry System) (weininger1988smiles). Duplicate entries were removed, and for pairs with varying PCEs due to fabrication differences, the highest reported value was retained to align with our model’s goal of predicting maximum achievable PCE. Data formats were standardized through manual inspection and corrections using the original literature. To ensure consistency, simplified alkyl chains in some datasets were restored to their authentic forms.
@@
To estimate the likelihood of organic photovoltaic (OPV) activity, we train a binary Random Forest (RF) classifier (breiman2001random). The dataset comprises 1,500 OPV-active molecules from the OPV2D dataset and 1,500 presumed-inactive molecules randomly selected from PubChem (kim2025pubchem). Each molecule is encoded using RDKit-derived molecular descriptors:
@@
The molecular encoder uses a three-stage GATv2-based hierarchy (brody2021attentive) to capture local and global interactions. Feature propagation at each layer is defined as:
@@
where, $\mathbf{z}_{i}^{(l)}$ denotes the feature vector of node $i$ at layer $l$, and $\mathcal{M}(i)$ represents its neighboring nodes. The operator $\bigoplus$ is an attention-based message aggregation function with 8 attention heads, and $\psi(\mathbf{z}_{j}^{(l)},\mathbf{e}_{ij})$ incorporates edge attributes $\mathbf{e}_{ij}$ from molecular graph, which encode bond properties such as type, aromaticity, and conjugation. Molecular graphs are derived from SMILES using RDKit (rdkit), with atoms as nodes and bonds as edges. An edge-conditioned dynamic attention mechanism allows the model to distinguish between different interatomic interactions, such as covalent bonds, and $\pi$$\pi$.
@@
The Material Generative Pretrained Transformer (MatGPT) extends the GPT-2 architecture (radford2019language)—a stack of masked transformer decoders—with targeted changes to improve molecular generation. Designed to prioritize both chemical validity and diversity, MatGPT learns from SMILES sequences to generate novel molecules. Our baseline implementation retains GPT’s core components: 8 transformer layers, each containing 8 multi-head self-attention heads, followed by position-wise feed-forward networks. The model processes sequences of 340 tokens using 256-dimensional embeddings.
Two key architectural modifications improve the model’s ability to capture molecular structural constraints and enhance generation diversity. (1) We adopt Rotary Positional Encoding (RoPE) (su2024roformer), which replaces absolute positional embeddings with rotational transformations, enabling implicit encoding of cyclic and relative patterns common in molecular graphs. (2) We introduce a Gated Linear Unit (GLU)-based feed-forward layer that better captures local and global feature interactions. Specifically, the hidden representation $x\in\mathbb{R}^{d}$ is transformed as follows:
@@
where $\odot$ denotes element-wise multiplication and $\sigma$ is the GELU (hendrycks2016gaussian) activation function. This formulation allows dynamic gating of features while maintaining computational efficiency. To further improve molecular generation, we implement a diversified sampling method that combines top-$p$ filtering (holtzman2019curious), and temperature annealing. At each generation step, the logits are adjusted using a dynamic temperature schedule $T(t)$, interpolating between $T_{\text{max}}$ and $T_{\text{min}}$ over time:
@@
where $\hat{y}_{\text{PCE}}$ is the predicted power conversion efficiency by P3, $P(\text{OPV}\mid\mathbf{x(u)})$ is the OPV likelihood calculated by OPVC, and $\mathbb{I}_{\text{valid}}$ is the validity indicator function. It is constrained by the Shockley-Queisser theoretical efficiency limit (shockley2018detailed). The KL divergence term provides statistical control over electronic properties:
@@
This framework incorporates experience replay (lin1992self), strategically reintegrating the top-performing 5% of historical molecules into subsequent training batches to reinforce successful chemical motifs. To maintain molecular diversity, a similarity threshold based on the Tanimoto index (>> 0.7) is applied to suppress redundant structures.
@@
We trained OPVC on a dataset consists of 1,500 positive examples from the OPV2D dataset and 1,500 negative samples randomly drawn from the PubChem database (kim2025pubchem), using 8:2 train-test split. The model was configured with 200 estimators, a maximum depth of 15, and class-balanced weighting. It achieved excellent performance, with a test accuracy of 99.16% and F1 score of 0.9916. The AUC reached 0.9985 on the test set, indicating strong generalization and robust classification performance.
@@
We pre-trained the graph encoder of MOE2 using three tasks: molecular mask reconstruction, computed HOMO-LUMO prediction, and literature-reported HOMO-LUMO prediction. For the first two tasks, we used over 51,000 non-fullerene acceptors (NFAs) from the reported dataset (Lopez201751k), and for the third task, we used the proposed OPV2D dataset. A 9:1 train-test split was employed for all tasks. The model achieved 99.997% accuracy on the reconstruction task and $R^{2}$ scores of 0.9866 and 0.9784 for the computed and literature-reported HOMO-LUMO prediction tasks, respectively.
To ensure a fair comparison in the baseline experiments, we also pre-trained other graph encoders using the same three tasks and datasets. Three architectures—GCNConv (kipf2016semi), GATv2Conv (brody2021attentive), and GINConv (xu2018powerful)—were trained under identical hyperparameter settings.
@@
After pretraining, we evaluated the learned graph embeddings against traditional Morgan fingerprints on the OPV2D dataset. For each sample, the embeddings were concatenated and used as input to a range of regression models, including Random Forest (RF) (breiman2001random), Support Vector Machine (SVM) (cortes1995support), Gradient Boosting (GB) (friedman2001greedy), XGBoost (chen2016xgboost), CatBoost (prokhorenkova2018catboost), LightGBM (LGBM) (ke2017lightgbm), Gaussian Processes (GP) (seeger2004gaussian), Histogram-based Gradient Boosting (HGB) (pedregosa2011scikit), and Natural Gradient Boosting (NGB) (duan2020ngboost), to predict power conversion efficiency (PCE). We used 5-fold cross-validation and the coefficient of determination ($R^{2}$) as the evaluation metric. Baseline results indicated that GAT-based embeddings, which incorporate edge weight information, consistently outperformed traditional approaches.
@@
GNN
embedding
GAT
embedding
GIN
embedding
@@
To evaluate the molecular generation capability of MatGPT, we adopted the Molecular Sets (MOSES) (polykovskiy2020molecular) benchmarking platform with protocol adaptations for organic photovoltaic (OPV) applications. The MOSES metrics were selected based on their ability to address three critical requirements for OPV material discovery: (1) Validity to ensure synthesizability of donor/acceptor candidates, (2) Novelty to avoid rediscovery of known low-efficiency materials, and (3) Diversity (IntDiv, IntDiv2) to explore broad chemical spaces for synergistic pairings. Besides vanilla GPT-2, we compared MatGPT against baseline models spanning diverse generative paradigms, including VAE (autoencoder-based) (gomez2018automatic), ORGAN (Generative Adversarial Networks) (guimaraes2017objective), a combinatorial generator (template-based) (polykovskiy2020molecular), and NGram (rule-based) (brown1992class).
@@
We selected PubChem (kim2025pubchem) for pretraining due to its broader chemical diversity compared to drug-focused datasets like ZINC Clean Leads (sterling2015zinc) and ChEMBL (gaulton2012chembl). Domain-specific preprocessing was applied to tailor the data for OPV materials, filtering SMILES strings to 15–300 characters and excluding ionic or multi-molecule entries. From the cleaned data, 5 million molecules were sampled to train MatGPT-Large, while MatGPT, GPT-2, and other baselines used a 1 million subset. An 8:2 split was used for training and reference sets. MatGPT and GPT-2 were trained for 30 epochs on the Causal Language Modeling task (radford2018improving) using a cosine annealing schedule (initial learning rate: 1e-4); other baselines followed MOSES defaults for fair comparison. Each model generated 30,000 molecules, evaluated with MOSES metrics. MatGPT used top-p sampling (p = 0.91) and temperature scaling (0.5–1.4) for diversity. All experiments ran on a single NVIDIA RTX 4090 GPU.
@@
We begin with the pre-trained MatGPT-Large model and fine-tune it on the OPV2D dataset, updating only the gated linear unit (GLU) layers while freezing all other parameters. This is followed by RL–based fine-tuning, where all model parameters remain frozen except for the GLU layers. During RL, we use the trained OPVC to estimate OPV activity, MOE2 for predicting HOMO and LUMO energy levels, and P3 for predicting power conversion efficiency (PCE). The loss function includes a reward scaling coefficient $\zeta_{1}=100$. The RL fine-tuning was run for 1,000 steps with a batch size of 256 and a learning rate of $1\times 10^{-4}$, completing in approximately 3 hours on a single NVIDIA RTX 4090 GPU.
@@
To validate the practical effectiveness of our framework, we selected representative candidates from two design tasks for quantum chemistry analysis. All DFT and TD-DFT calculations were performed using the Gaussian 16 and GaussView 6 software package (g16; gv6). In organic photovoltaics, two factors are critical for high performance: (1) Energy Level Alignment and (2) Spectral Complementarity. Proper alignment of the donor-acceptor frontier molecular orbitals (HOMO/LUMO) is essential to create a driving force for separating light-induced excitons into free charges. Concurrently, the donor and acceptor should absorb different regions of the solar spectrum, working in concert to achieve panchromatic light harvesting and maximize current generation.
@@
Our framework demonstrated success in optimizing both factors, as illustrated in Figure 6. The figure displays the calculated monomer absorption spectra (Monomer, Calc.) for our representative generated molecules and their estimated solid-state film spectra (Film, Est.). Guided by experimental literature (Y6; Y6-2; PTB7), which typically reports a 100-200 nm redshift from solution to solid-state films due to increased molecular planarity and strong intermolecular $\pi-\pi$ interactions, we selected a representative shift of 150 nm for our estimation. For the PTB7-Th pair, our generated acceptor extends the absorption deep into the near-infrared region (700-900 nm), perfectly complementing the donor’s visible light absorption. Conversely, for the Y6 pair, our generated donor effectively fills the visible spectrum gap (400-650 nm) left by the NIR-absorbing Y6. In addition to this ideal spectral complementarity, quantum chemistry calculations confirmed that both pairs possess well-aligned frontier molecular orbital energies conducive to efficient exciton dissociation. This dual validation of optical and electronic properties underscores our framework’s ability to discover novel, high-potential donor-acceptor partners.

### v2

@@ Abstract
Codehttps://github.com/HouHei0416/CycleChemist
Datasetshttps://github.com/sunyrain/OPV2D
@@ Introduction
Current machine learning models used for predicting the power conversion efficiency (PCE) of organic photovoltaic (OPV) materials remain relatively simplistic. Limitations in both algorithm design and dataset sizes have resulted in insufficient predictive accuracy and generalizability. For example, a comprehensive survey of various embedding methods and machine learning models was conducted using 558 data points, with tree-based algorithms achieving the highest predictive correlation (approximately 0.61) (Seifrid et al. 2024). However, such a small dataset may not adequately represent mainstream OPV material systems, thereby significantly affecting generalizability. In another study, a larger dataset was compiled and a random forest model was trained, reporting an $R^{2}$ value of 0.71 (Miyake and Saeki 2021). Nonetheless, the reliance on experimental data as model input renders this approach impractical for large-scale screening studies. In (Zhang et al. 2025b), training was performed on 1,060 donor–acceptor pairs using various graph neural network algorithms, with the Graph Attention Network (GAT) achieving the best result (reported $r$ value of 0.74). Additionally, the model presented in (Das and Mondal 2024) was trained on 300 donor–acceptor pairs and involved detailed quantum mechanical calculations for each molecule during dataset preparation, introducing new quantum mechanical descriptors. On this small dataset, the gradient boosting (GB) model achieved the best performance with a reported $r$ value of 0.92. However, the complexity of the quantum mechanical (QM) calculations limits the model’s potential for large-scale screening. Similarly, another study utilizing only 100 data points incorporated numerous QM descriptors and reported an $r$ value of 0.918 (Khatua et al. 2024). In summary, these methods are either trained on relatively small datasets or require complex QM calculations or additional experimental parameters, all of which limit their applicability for large-scale screening and may lead to performance that is somewhat contingent on chance.
@@
Previous research has predominantly focused on predicting the properties of photovoltaic materials, and in recent years, some scholars have begun to explore donor–acceptor design based on machine learning. According to the literature, methods for exploring high-efficiency OPV materials mainly include strategies such as disassembling and recombining existing molecular structures, as well as employing variational autoencoders (VAE) and genetic algorithms (GA) (Greenstein and Hutchison 2023; Sun et al. 2024; Zhang et al. 2025a). These approaches largely rely on optimizing known molecules, thereby limiting their capacity to generate and explore entirely novel molecules. Some studies do not involve molecular design at all but instead exhaustively enumerate possible donor–acceptor combinations, exploring potential high-efficiency pairs within a predefined design space (Khatua et al. 2024). Moreover, most related studies employ datasets comprising only 100 to 1,500 samples (Miyake and Saeki 2021; Wu et al. 2020; Greenstein and Hutchison 2023), covering relatively narrow material systems, which consequently restricts the diversity of molecular fragments that the models can explore.
@@
Based on previous works (Miyake and Saeki 2021; Wu et al. 2020), we integrated and curated multiple organic photovoltaic datasets to construct the Organic Photovoltaic Donor-Acceptor Dataset (OPV2D), comprising approximately 2,000 donor–acceptor pairs—one of the largest OPV datasets to date. Molecules are represented using SMILES (Simplified Molecular Input Line Entry System) (Weininger 1988). Duplicate entries were removed, and for pairs with varying PCEs due to fabrication differences, the highest reported value was retained to align with our model’s goal of predicting maximum achievable PCE. Data formats were standardized through manual inspection and corrections using the original literature. To ensure consistency, simplified alkyl chains in some datasets were restored to their authentic forms.
@@
To estimate the likelihood of organic photovoltaic (OPV) activity, we train a binary Random Forest (RF) classifier (Breiman 2001). The dataset comprises 1,500 OPV-active molecules from the OPV2D dataset and 1,500 presumed-inactive molecules randomly selected from PubChem (Kim et al. 2025). Each molecule is encoded using RDKit-derived molecular descriptors:
@@
The molecular encoder uses a three-stage GATv2-based hierarchy (Brody et al. 2021) to capture local and global interactions. Feature propagation at each layer is defined as:
@@
where, $\mathbf{z}_{i}^{(l)}$ denotes the feature vector of node $i$ at layer $l$, and $\mathcal{M}(i)$ represents its neighboring nodes. The operator $\bigoplus$ is an attention-based message aggregation function with 8 attention heads, and $\psi(\mathbf{z}_{j}^{(l)},\mathbf{e}_{ij})$ incorporates edge attributes $\mathbf{e}_{ij}$ from molecular graph, which encode bond properties such as type, aromaticity, and conjugation. Molecular graphs are derived from SMILES using RDKit (RDKit, online 2013), with atoms as nodes and bonds as edges. An edge-conditioned dynamic attention mechanism allows the model to distinguish between different interatomic interactions, such as covalent bonds, and $\pi$$\pi$.
@@
The Material Generative Pretrained Transformer (MatGPT) extends the GPT-2 architecture (Radford et al. 2019)—a stack of masked transformer decoders—with targeted changes to improve molecular generation. Designed to prioritize both chemical validity and diversity, MatGPT learns from SMILES sequences to generate novel molecules. Our baseline implementation retains GPT’s core components: 8 transformer layers, each containing 8 multi-head self-attention heads, followed by position-wise feed-forward networks. The model processes sequences of 340 tokens using 256-dimensional embeddings.
Two key architectural modifications improve the model’s ability to capture molecular structural constraints and enhance generation diversity. (1) We adopt Rotary Positional Encoding (RoPE) (Su et al. 2024), which replaces absolute positional embeddings with rotational transformations, enabling implicit encoding of cyclic and relative patterns common in molecular graphs. (2) We introduce a Gated Linear Unit (GLU)-based feed-forward layer that better captures local and global feature interactions. Specifically, the hidden representation $x\in\mathbb{R}^{d}$ is transformed as follows:
@@
where $\odot$ denotes element-wise multiplication and $\sigma$ is the GELU (Hendrycks and Gimpel 2016) activation function. This formulation allows dynamic gating of features while maintaining computational efficiency. To further improve molecular generation, we implement a diversified sampling method that combines top-$p$ filtering (Holtzman et al. 2019), and temperature annealing. At each generation step, the logits are adjusted using a dynamic temperature schedule $T(t)$, interpolating between $T_{\text{max}}$ and $T_{\text{min}}$ over time:
@@
where $\hat{y}_{\text{PCE}}$ is the predicted power conversion efficiency by P3, $P(\text{OPV}\mid\mathbf{x(u)})$ is the OPV likelihood calculated by OPVC, and $\mathbb{I}_{\text{valid}}$ is the validity indicator function. It is constrained by the Shockley-Queisser theoretical efficiency limit (Shockley and Queisser 2018). The KL divergence term provides statistical control over electronic properties:
@@
This framework incorporates experience replay (Lin 1992), strategically reintegrating the top-performing 5% of historical molecules into subsequent training batches to reinforce successful chemical motifs. To maintain molecular diversity, a similarity threshold based on the Tanimoto index (>> 0.7) is applied to suppress redundant structures.
@@
We trained OPVC on a dataset consists of 1,500 positive examples from the OPV2D dataset and 1,500 negative samples randomly drawn from the PubChem database (Kim et al. 2025), using 8:2 train-test split. The model was configured with 200 estimators, a maximum depth of 15, and class-balanced weighting. It achieved excellent performance, with a test accuracy of 99.16% and F1 score of 0.9916. The AUC reached 0.9985 on the test set, indicating strong generalization and robust classification performance.
@@
We pre-trained the graph encoder of MOE2 using three tasks: molecular mask reconstruction, computed HOMO-LUMO prediction, and literature-reported HOMO-LUMO prediction. For the first two tasks, we used over 51,000 non-fullerene acceptors (NFAs) from the reported dataset (Lopez et al. 2017), and for the third task, we used the proposed OPV2D dataset. A 9:1 train-test split was employed for all tasks. The model achieved 99.997% accuracy on the reconstruction task and $R^{2}$ scores of 0.9866 and 0.9784 for the computed and literature-reported HOMO-LUMO prediction tasks, respectively.
To ensure a fair comparison in the baseline experiments, we also pre-trained other graph encoders using the same three tasks and datasets. Three architectures—GCNConv (Kipf 2016), GATv2Conv (Brody et al. 2021), and GINConv (Xu et al. 2018)—were trained under identical hyperparameter settings.
@@
After pretraining, we evaluated the learned graph embeddings against traditional Morgan fingerprints on the OPV2D dataset. For each sample, the embeddings were concatenated and used as input to a range of regression models, including Random Forest (RF) (Breiman 2001), Support Vector Machine (SVM) (Cortes and Vapnik 1995), Gradient Boosting (GB) (Friedman 2001), XGBoost (Chen and Guestrin 2016), CatBoost (Prokhorenkova et al. 2018), LightGBM (LGBM) (Ke et al. 2017), Gaussian Processes (GP) (Seeger 2004), Histogram-based Gradient Boosting (HGB) (Pedregosa et al. 2011), and Natural Gradient Boosting (NGB) (Duan et al. 2020), to predict power conversion efficiency (PCE). We used 5-fold cross-validation and the coefficient of determination ($R^{2}$) as the evaluation metric. Baseline results indicated that GAT-based embeddings, which incorporate edge weight information, consistently outperformed traditional approaches.
@@
GNN embedding GAT embedding GIN embedding
@@
To evaluate the molecular generation capability of MatGPT, we adopted the Molecular Sets (MOSES) (Polykovskiy et al. 2020) benchmarking platform with protocol adaptations for organic photovoltaic (OPV) applications. The MOSES metrics were selected based on their ability to address three critical requirements for OPV material discovery: (1) Validity to ensure synthesizability of donor/acceptor candidates, (2) Novelty to avoid rediscovery of known low-efficiency materials, and (3) Diversity (IntDiv, IntDiv2) to explore broad chemical spaces for synergistic pairings. Besides vanilla GPT-2, we compared MatGPT against baseline models spanning diverse generative paradigms, including VAE (autoencoder-based) (Gómez-Bombarelli et al. 2018), ORGAN (Generative Adversarial Networks) (Guimaraes et al. 2017), a combinatorial generator (template-based) (Polykovskiy et al. 2020), and NGram (rule-based) (Brown et al. 1992).
@@
We selected PubChem (Kim et al. 2025) for pretraining due to its broader chemical diversity compared to drug-focused datasets like ZINC Clean Leads (Sterling and Irwin 2015) and ChEMBL (Gaulton et al. 2012). Domain-specific preprocessing was applied to tailor the data for OPV materials, filtering SMILES strings to 15–300 characters and excluding ionic or multi-molecule entries. From the cleaned data, 5 million molecules were sampled to train MatGPT-Large, while MatGPT, GPT-2, and other baselines used a 1 million subset. An 8:2 split was used for training and reference sets. MatGPT and GPT-2 were trained for 30 epochs on the Causal Language Modeling task (Radford et al. 2018) using a cosine annealing schedule (initial learning rate: 1e-4); other baselines followed MOSES defaults for fair comparison. Each model generated 30,000 molecules, evaluated with MOSES metrics. MatGPT used top-p sampling (p = 0.91) and temperature scaling (0.5–1.4) for diversity. All experiments ran on a single NVIDIA RTX 4090 GPU.
@@
We begin with the pre-trained MatGPT-Large model and fine-tune it on the OPV2D dataset, updating only the gated linear unit (GLU) layers while freezing all other parameters. This is followed by RL–based fine-tuning, where all model parameters remain frozen except for the GLU layers. During RL, we use the trained OPVC to estimate OPV activity, MOE2 for predicting HOMO and LUMO energy levels, and P3 for predicting power conversion efficiency (PCE). The loss function includes a reward scaling coefficient $\zeta_{1}=100$. The RL fine-tuning was run for 500 steps with a batch size of 128 and a learning rate of $1\times 10^{-4}$, completing in approximately 3 hours on a single NVIDIA RTX 4090 GPU.
@@
To validate the practical effectiveness of our framework, we selected representative candidates from two design tasks for quantum chemistry analysis. All DFT and TD-DFT calculations were performed using the Gaussian 16 and GaussView 6 software package (Frisch et al. 2016; Dennington et al. 2019). In organic photovoltaics, two factors are critical for high performance: (1) Energy Level Alignment and (2) Spectral Complementarity. Proper alignment of the donor-acceptor frontier molecular orbitals (HOMO/LUMO) is essential to create a driving force for separating light-induced excitons into free charges. Concurrently, the donor and acceptor should absorb different regions of the solar spectrum, working in concert to achieve panchromatic light harvesting and maximize current generation.
@@
Our framework demonstrated success in optimizing both factors, as illustrated in Figure 6. The figure displays the calculated monomer absorption spectra (Monomer, Calc.) for our representative generated molecules and their estimated solid-state film spectra (Film, Est.). Guided by experimental literature (Yuan et al. 2019; P Christopholi et al. 2024; Li et al. 2021), which typically reports a 100-200 nm redshift from solution to solid-state films due to increased molecular planarity and strong intermolecular $\pi-\pi$ interactions, we selected a representative shift of 150 nm for our estimation. For the PTB7-Th pair, our generated acceptor extends the absorption deep into the near-infrared region (700-900 nm), perfectly complementing the donor’s visible light absorption. Conversely, for the Y6 pair, our generated donor effectively fills the visible spectrum gap (400-650 nm) left by the NIR-absorbing Y6. In addition to this ideal spectral complementarity, quantum chemistry calculations confirmed that both pairs possess well-aligned frontier molecular orbital energies conducive to efficient exciton dissociation. This dual validation of optical and electronic properties underscores our framework’s ability to discover novel, high-potential donor-acceptor partners.